# Retrograde predominance of small saturnian moons reiterates a recent retrograde collisional disruption

### v1

@@ 4.2 Gallic Group
The Gallic group only gained three new members, which makes it now 9 moons in total (see Figure 2). Of the six previously known Gallic group members, all but S/2004 S 24 have similar $a$. This clustering, which we will refer to as the Albiorix subgroup, also includes all three of the new members. Interestingly, Albiorix has the lowest $a$ of all known Gallic members. Perhaps the Albiorix subgroup, or even the Gallic group as a whole, was created with a head-on collision resulting in the largest fragment, Albiorix, having an orbit close to that of the parent body. All the smaller fragments would get a $\Delta v$ boost relative to the parent body in roughly the direction of motion, thus producing larger $a$. Given that there are significantly more retrograde irregular moons compared to direct ones, a head-on collision is a plausible scenario.
One of the new Gallic moons, S/2006 S 12, has reduced the size of the semimajor axis gap between the Albiorix subgroup and S/2004 S 24 but this gap still remains larger than the pericentre of S/2004 S 24. The disconnect of S/2004 S 24 from the rest of the group, in terms of having a significantly larger $a$ and being the only Gallic member whose perihelion is not lower than the semi-major axis of Albiorix itself, means that the question still remains whether S/2004 S 24 comes from the same parent body as the rest of the Gallic group. The apocenters of all the Gallic members cross the $a$ of S/2004 S 24, so perhaps there was a secondary collision in the Gallic group where an Albiorix subgroup member was impacted while near apocenter and S/2004 S 24 is the largest fragment of that collision. More Gallic group members will need to be discovered to better understand the origin of S/2004 S 24.
@@ 4.3 Inuit Group
The Inuit group has more than doubled their number of known members, from 6 to 13. With the addition of these new Inuit moons, some structure within the group has become clear (see Figure 2). All of the new Inuit moons have semi-major axes close to that of a previously known moon, producing two subgroups which were both briefly mentioned by Sheppard et al. (2023).
@@ 4.4 Norse Group
Gladman et al. (2001) initially suggested that Skathi (S/2000 S 8) should be separated from the rest of the so-called Phoebe subgroup due to its significantly different orbital inclination; in the end all retrograde saturnians have been receiving norse-themed names. Even the postulated ‘Phoebe group’ was remarked to be unusual, in that Phoebe is very much larger than the other 4 discoveries from 2000: Ymir, Thrymr, Suttungr (which have very Phoebe-like inclinations but much larger semimajor axes) and Muldilfari, which has an inclination offset from Phoebe much larger than the spread seen in the two direct inclination groups).
Like the previously known moons, the vast majority of the new moons belong to the Norse group. Of the 64 new moons, 54 are Norse group members. This is a slightly larger fraction than that of previously known moons (46/58). The retrograde population has now exploded in number into a very complicated distribution that is far more difficult to interpret than the tightly-clustered direct groups. At first glance, the rose plot of the retrograde moons (Figure 3) does not cleanly manifest much structure in the orbital distribution. Looking at a cumulative inclination distribution, however, it appears that the Norse group can be broken up by using inclination cuts tied to subtle features (top panel of Figure 4).
Figure 3: Rose diagram of the retrograde saturnian irregulars. See Fig. 2’s caption. Here, red represents moons belonging to the Phoebe subgroup, magenta represents the Mundilfari subgroup, blue represents the Kari subgroup, and cyan represents low-i ($<$151) Norse members. Two Phoebe subgroup members that may be outlying members of the Mundilfari subgroup are indicated with an additional inset magenta dot.
Figure 4: Inverse cumulative inclination distribution of the full Norse/retrograde population (top panel) and then split up into large ($>4$ km) and small ($<4$ km) members (bottom panel). We divide the Norse group up into three sub-groups with red dashed lines marking the boundaries. The short red dotted line in each panel indicates a dearth of moons within the Mundilfari subgroup (discussed in subsection 5.1).
@@ 4.4.1 Phoebe Subgroup
Ashton et al. (2021) attempted to subdivide the Norse group by defining the Phoebe subgroup as moons with inclinations within 3 of Phoebe’s value. From the top panel of Figure 4 a clear absence of moons (relative to surrounding inclinations) can be seen at 172.5, suggesting there is a division between Phoebe-like inclinations and the rest. This division can be seen more clearly when looking at just the large (D $>$ 4 km) Norse members (bottom panel of Figure 4). Here we will use the criteria $i>172.5^{\circ}$ to define the Phoebe subgroup (although, currently, we get the same Phoebe subgroup members if we use the criteria from Ashton et al. (2021)). The recent batch of new discoveries has added 10 more members to the Phoebe subgroup for a total of 21 (see footnote 6 for the full list).
The saturnocentric distance variation of a large fraction of the Phoebe subgroup members do not overlap with that of Phoebe (i.e. the pericenter of some members is larger than the apocenter of Phoebe). As such, a scenario of the subgroup being created by a single collision off Phoebe seems unlikely. There is a small clustering in Phoebe subgroup members around 20×$\times 10^{6}$ $\simeq$0.13 au which may be a good candidate for a small collisional subfamily. Furthermore, there are four moons, Angrboda, Alvaldi, S/2019 S 13, and S/2004 S 46, which have similar $a$, and almost the same $i$, 177.2–177.4. This close clustering in $i$, and the fact that they are 1.6 away from the rest of the Phoebe subgroup, could be an indication of another small collisional subfamily. Curiously, Ymir is still quite isolated in $a-i$ space. Lastly, the lowest-$a$ members of the Phoebe subgroup, S/2006 S 9, S/2006 S 20, S/2019 S 2, and S/2007 S 2, could be fragments of a crating event off Phoebe.
@@ 4.4.2 Kari Subgroup and low inclination Norse moons
At the low-$i$ end of the Norse group (closest to $i=90^{\circ}$) the cumulative distribution (top panel of Figure 4) flattens out when going to smaller $i$. This flattening occurs at two $i$’s, 157 and 151. This suggests that these $i$’s might be good locations for having subgroup boundaries. As such, we define the Kari subgroup (named after the largest member) as the moons with ${}^{\circ}<$ $i<$ 157. We could easily put the lower dividing line at $i$ = 152, but since S/2019 S 19 has an $a$ that is similar to a majority of the members in the Kari subgroup we suggest it should be included in the subgroup. The tightly packed nature of the Kari subgroup, without the 2-3 lowest $a$ members, makes it a good candidate for a collisional family
Of the five Norse moons with the lowest $i$, which were the only moons not placed in any subgroups, Narvi and S/2019 S 11 probably share a common origin. Skathi and Hyrrokkin could potentially be related (although there is a large separation between their $a$’s), and Bestla is by itself. We note that this population is approaching $i=90^{\circ}$ closely enough that the Kozai effect may begin to operate, like that observed for three Inuit moons by Jacobson et al. (2022), with the potential for long-term instability of the moons Gladman et al. (2001) setting a boundary for this group.
@@ 4.4.3 Mundilfari Subgroup
Using the inclination distribution is a good method to define Norse subgroups as collisional fragments will clump together in $i$. Although, having just a single $i$ as the boundary between groups will most likely result in moons being put into sub-groups that do not belong to. Two good examples of this are discussed in section 5.
@@ 5 Size Distribution
Although the inclination study was done first, and the boundaries of these divisions are obviously not very strong, we later discovered that the splits as they made them resulted in a dramatically different size distribution for the Mundilfari subgroup when compared to all other saturnians. When one splits up the saturnian irregular size distribution into different groups/subgroups (Figure 5) one can see that the Mundilfari subgroup is significantly steeper than the other clusters over a large range of diameters. As shown by one of the reference slopes (dotted line) in Figure 5 the Mundilfari subgroup has a fairly consistent differential power law index, $q$, of approximately 6 down until at least $D\sim$ 3 km, at which point incompleteness causes the distribution to flatten out.
@@
The other two Norse subgroups, Phoebe and Kari, generally have much less steep slopes in the same size range (3–6 km), similar to that of collisional equilibrium, $q=3.5$ (Dohnanyi, 1969). Even shallower still are the two direct groups, Gallic and Inuit, which both have a shallow slopes of $q\sim$ 2 throughout.
Both the Phoebe and Kari subgroups have a narrow size range around 4 km where the size distributions have similar steepness to that of the Mundilfari subgroup. For the Kari subgroup this might be due to small-number statistics. More Kari members need to be found to get a better understanding of the sub-group’s size distribution. As mentioned in subsubsection 4.4.3, there is likely cross-contamination between what we have defined as subgroups. In the size range where the Phoebe subgroup has a very steep slope (3.7–4.7 km), 3 out of the 8 members have $i$ within 1 of the Mundilfari subgroup (i.e. 172.5 $<i<$ 173.5). One of the three moons is S/2006 S 20 which, as mentioned in subsubsection 4.4.1 is likely a collisional fragment off of Phoebe. The other two moons, Griep and Sat LVIII, could easily belong to the Mundilfari subgroup, as they have similar $a$ to Mundilfari members just on the other side of the inclination divide (see Figure 3). If both of these moons are indeed members of the Mundilfari subgroup and not the Phoebe subgroup, the slope of the size distribution at around 4 km would be significantly shallower for the Phoebe subgroup.
Ashton et al. (2021) found that the saturnian irregulars, as a whole, had $q=4.9$ from $D=3$ to 4 km and postulated that this steep slope was due to a recent collision (in the last couple of Gyr). They were unable to determine which members were likely involved in the collision as almost all of the moons that they detected in the size range did not have known orbits at the time. However, they did suggest that it was likely located in the retrograde population. The new saturnian irregulars discovered show that the Mundilfari subgroup alone is essentially the cause of the steep size distribution found by Ashton et al. (2021). Thus, we believe that the Mundilfari subgroup was created by the recent collision hypothesised by Ashton et al. (2021). If this is true then the Mundilfari collisional family is significantly more dispersed in $a$ and $i$ than the other potential collisional families seen among the irregular moons.
@@ 5.1 Splitting up the Mundilfari Subgroup
Perhaps this steep slope is only present in a subset of the Mundilfari subgroup. If the subgroup is split into two by a potential division at $i$ = 163, as discussed in subsubsection 4.4.3, the slope of the size distributions of the two regions are very similar to that of the Mundilfari subgroup as a whole (see left panel in Figure 6). Additionally, we split the Mundilfari subgroup into two subsamples two different ways using cuts of $a=2.1\times 10^{7}$ km or $e=0.25$. The values of the $a$ and $e$ cuts were chosen to split the Mundilfari subgroup into roughly equal numbers. Like with the $i$ split, the size distribution of the high/low $e$/$a$ all have very high slopes, like the Mundilfari subgroup as a whole (middle and right panels in Figure 6, respectively); the $e$ $<$ 0.25 sample has the shallowest slope with $q\sim 5$. We therefore believe that the steep slope present in all subdivisions justifies our choice to keep the Mundilfari subgroup intact.
@@ 5.2 Observational Bias
A caveat that must be taken into consideration is that is sample has not been debaised. Even though a majority of saturnian irregulars lie in the two offset fields (Fig. 1) at any one instant, and the two fields have been observed many times over multiple oppositions, this survey is still biased towards moons with certain orbits. The lack of coverage north and south of Saturn (see Figure 1) results in moons with $i$ closer to 90 being harder to detect compared to moons with $i$ closer to 0 or 180. This is exemplified by the fact that we never imaged Bestla, the lowest $i$ Norse group member. Another bias, as mentioned in Ashton et al. (2022), is the glare from Saturn, making it harder to observe moons with smaller $a$ or pericentre. Given these two biases, the Inuit and Gallic subgroups and the low-$i$ Norse moons ($i<$ 151) are likely underrepresented in the current sample, and thus have size distributions that are likely more steep in reality than what is seen in Figure 5. The Norse subgroups, on the other hand, are less effected by such observational biases, with the possible exceptions of the low-$a$ part of the Phoebe subgroup and, to a lesser extent, the Kari subgroup. Therefore, we believe that observational biases in our survey should not influence our analysis or conclusions regarding the Mundilfari subgroup.
6 Mundilfari Subgroup; Collision Family?
6.1 Simulated collision
To test the feasibility of a single collision creating the Mundilfari subgroup we calculate the orbital extent of a debris field given a certain ejection speed. For the orbit of the progenitor moon, we used the approximate center of the Mundilfari subgroup for $i$, $a$, and $e$, with values of 165, $1.95\times 10^{7}$ km (0.126 au), and 0.28 respectively. We varied the longitude of ascending node, $\Omega$, argument of periapsis, $\omega$, and true anomaly, $f$, to find an ejecta field that best matched the orbits of the Mundilfari subgroup.

### v2

@@ 4.2 Gallic Group
The Gallic group only gained three new members, which makes it now 9 moons in total (see Figure 2). Of the six previously known Gallic group members, all but S/2004 S 24 have similar $a$. This clustering, which we will refer to as the Albiorix subgroup, also includes all three of the new members. Interestingly, Albiorix has the lowest $a$ of all known Gallic members. Perhaps the Albiorix subgroup, or even the Gallic group as a whole, was created via a cratering event on the leading hemisphere of Albiorix, with this still (mostly) intact parent body remaining close to its pre-collision orbit. If the crater location was near the apex point of Albiorix’s velocity, all the fragments would be ejected in the direction of the leading hemisphere; their relative $\Delta v$ boost would have some component in the direction of motion of the parent body, producing larger $a$ (see subsection 6.1). Such a leading-hemisphere impact event is greatly favoured when the projectile is moving in the opposite orbital sense from the target, and naturally results in a higher relative impact speed (about twice the orbital speed for a direct encounter), allowing retrograde projectiles to provide a much larger kinetic energy to the impact. Given that there are significantly more retrograde irregular moons compared to direct ones, a leading-hemisphere collision is a plausible scenario.
One of the new Gallic moons, S/2006 S 12, has reduced the size of the semi-major axis gap between the Albiorix subgroup and S/2004 S 24 but this gap still remains larger than the pericentre of S/2004 S 24. The disconnect of S/2004 S 24 from the rest of the group, in terms of having a significantly larger $a$ and being the only Gallic member whose perihelion is not lower than the semi-major axis of Albiorix itself, means that the question still remains whether S/2004 S 24 comes from the same parent body as the rest of the Gallic group. We explore this further in subsection 6.1.
@@ 4.3 Inuit Group
The Inuit group has more than doubled their number of known members, from 6 to 13. With the addition of these new Inuit moons, some structure within the group has become clear (see Figure 2). All of the new Inuit moons have semi-major axes close to that of a previously known moon, producing two subgroups which were both mentioned by Sheppard et al. (2023).
@@ 4.4 Norse Group
Gladman et al. (2001) initially suggested that Skathi (S/2000 S 8) should be kept separate from the rest of the so-called Phoebe subgroup due to its significantly different orbital inclination; in the end, all retrograde saturnians have been receiving norse-themed names. Even the postulated ‘Phoebe group’ was remarked to be unusual, in that Phoebe is very much larger than the other 4 discoveries from 2000: Ymir, Thrymr, Suttungr (which have Phoebe-like inclinations but much larger semi-major axes) and Mundilfari, which has an inclination offset from Phoebe much larger than the spread seen in the two direct inclination groups).
Like the previously known moons, the vast majority of the new moons belong to the Norse/retrograde group. Of the 64 new moons, 54 are Norse group members. This is a slightly larger fraction than that of previously known moons (46/58). The retrograde population has now exploded in number into a very complicated distribution that is far more difficult to interpret than the tightly-clustered direct groups. At first glance, the rose plot of the retrograde moons (Figure 3) does not cleanly manifest much structure in the orbital distribution. Looking at a cumulative inclination distribution, however, it appears that the Norse group can be broken up by using inclination cuts tied to subtle features (top panel of Figure 4).
Figure 3: Rose diagram of the retrograde saturnian irregulars. See Fig. 2’s caption. Here, red represents moons belonging to the Phoebe subgroup, magenta represents the Mundilfari subgroup, blue represents the Kari subgroup, and cyan represents low-i (<<151) Norse members. Mundilfari is represented by a square and is a darker shade of purple, in order to distinguish the moon from the rest of the subgroup named after it. Two Phoebe subgroup members that may be outlying members of the Mundilfari subgroup are indicated with an additional inset magenta dot.
Figure 4: Inverse cumulative inclination distribution of the full Norse/retrograde population (top panel) and then split up into large (>4>4 km) and small (<4<4 km) members (bottom panel). We divide the Norse group up into three sub-groups with red dashed lines marking the boundaries. There are two sets of short red dotted lines in each panel, one at 176.5 indicating a potential further subdivision within the Phoebe subgroup (see subsubsection 4.4.1) and one at 163 indicating a minor dearth of moons within the Mundilfari subgroup (see subsection 5.1).
@@ 4.4.1 Phoebe Subgroup
Ashton et al. (2021) attempted to subdivide the Norse group by defining the Phoebe subgroup as moons with inclinations within 3 of Phoebe’s value. From the top panel of Figure 4 a clear absence of moons (relative to surrounding inclinations) can be seen at 172, suggesting there is a division between Phoebe-like inclinations and the rest. This division can be seen more clearly when looking at just the large (D >> 4 km) Norse members (bottom panel of Figure 4). Here we will use the criteria $i>172^{\circ}$ to define the Phoebe subgroup (although, currently, we get the same Phoebe subgroup members if we use the criteria from Ashton et al. (2021)). The recent batch of new discoveries has added 10 more members to the Phoebe subgroup for a total of 21 (see footnote 7 for the full list).
[…]
@@ 4.4.2 Kari Subgroup and low inclination Norse moons
[…]
Of the five Norse moons with the lowest $i$, which were the only moons not placed in any subgroups, Skathi and Hyrrokkin could potentially be related (although there is a large separation between their $a$’s). Narvi, S/2019 S 11, and Bestla have similar $a$’s but are a bit spread in $i$. We note that this population is approaching $i=90^{\circ}$ closely enough that the Kozai effect may begin to operate, like that observed for three Inuit moons by Jacobson et al. (2022), with the potential for long-term instability of the moons Gladman et al. (2001) setting a boundary for this group.
@@ 4.4.3 Mundilfari Subgroup
Using the inclination distribution is a good method to define Norse subgroups as collisional fragments will clump together in $i$. Although, having just a single $i$ as the boundary between groups will most likely result in moons being put into sub-groups that they do not belong to. Two good examples of this are discussed in section 5.
@@ 5 Size Distribution
Although the inclination study was done first, and the boundaries of these divisions are obviously not very strong, we later discovered that the splits as we made them resulted in a dramatically different size distribution for the Mundilfari subgroup when compared to all other saturnians. When one splits up the saturnian irregular size distribution into different groups/subgroups (Figure 5) one can see that the Mundilfari subgroup is significantly steeper than the other clusters over a large range of diameters. As shown by one of the reference slopes (dotted line) in Figure 5 the Mundilfari subgroup has a fairly consistent differential power law index, $q$, of approximately 6 down until at least $D\sim$ 3 km, at which point incompleteness causes the distribution to flatten out.
@@
The other two Norse subgroups, Phoebe and Kari, generally have much less steep slopes in the same size range (3–6 km), similar to that of collisional equilibrium, $q=3.5$ (Dohnanyi, 1969). Even shallower still are the two direct groups, Gallic and Inuit, which both have shallow slopes of $q\sim$ 2 throughout.
Both the Phoebe and Kari subgroups have a narrow size range around 4 km where the size distributions have similar steepness to that of the Mundilfari subgroup. As mentioned in subsubsection 4.4.3, there is likely cross-contamination between what we have defined as subgroups. In the size range where the Phoebe subgroup has a very steep slope (3.7–4.7 km), 4 out of the 8 members have $i$ within 2 of the Mundilfari subgroup (i.e. 172 <i<<i< 174). One of the four moons is S/2006 S 20 which, as mentioned in subsubsection 4.4.1 is likely a collisional fragment off of Phoebe. Another is S/2007 S 3, with an $a$ of 0.1311 au, is at the outer edge of a large cluster within the Phoebe subgroup, and is thus more likely to be associated with this cluster and not to the Mundilfari subgroup. The other two moons, S/2007 S 2 and Sat LVIII, could easily belong to the Mundilfari subgroup, as they are fairly isolated in $a$-$i$ (see Figure 3). Additionally, Sat LVIII has a similar $a$ to Mundilfari members just on the other side of the inclination divide. If both of these moons are indeed members of the Mundilfari subgroup and not the Phoebe subgroup, the slope of the size distribution at around 4 km would be significantly shallower for the Phoebe subgroup. A similar argument could be made for the Kari subgroup. After Kari, the next three largest members, S/2006 S 3, S/2006 S 1, and Farbauti, are all within 1 of the Mundilfari boundary and all have $a$’s lower than the main Kari subgroup cluster. Additionally, small-number statistics might play a part.
Ashton et al. (2021) found that the saturnian irregulars, as a whole, had $q=4.9$ from $D=3$ to 4 km and postulated that this steep slope was due to a recent collision (in the last Gyr). Catastrophic collisions (and cratering events) are known to produce steep size distrbutions (see Durda et al. (2007) for example). They were unable to determine which members were likely involved in the collision as almost all of the moons that they detected in the size range did not have known orbits at the time. However, they did suggest that it was likely located in the retrograde population. The new saturnian irregulars discovered show that the Mundilfari subgroup alone is essentially the cause of the steep size distribution found by Ashton et al. (2021). Thus, we believe that it is the Mundilfari subgroup that was created by the recent collision hypothesised by Ashton et al. (2021). If this is true then the Mundilfari collisional family is significantly more dispersed in $a$ and $i$ than the other potential collisional families seen among the irregular moons.
@@ 5.1 Splitting up the Mundilfari Subgroup
We explore the idea that perhaps this steep slope is only present in a subset of the Mundilfari subgroup. If the subgroup is split into two by a potential division at $i$ = 163, as discussed in subsubsection 4.4.3, the slope of the size distributions of the two regions are very similar to that of the Mundilfari subgroup as a whole (see left panel in Figure 6). Additionally, we split the Mundilfari subgroup into two subsamples two different ways using cuts of $a$ = 0.14 au or $e=0.25$. The values of the $a$ and $e$ cuts were chosen to split the Mundilfari subgroup into roughly equal numbers. Like with the $i$ split, the size distribution of the high/low ee/$a$ all have very high slopes, like the Mundilfari subgroup as a whole (middle and right panels in Figure 6, respectively); the $e$ << 0.25 sample has the shallowest slope with $q\sim 5$. We therefore believe that the steep slope present in all subdivisions justifies our choice to keep the Mundilfari subgroup intact.
@@ 5.2 Observational Bias
A caveat that must be taken into consideration is that is sample has not been debaised. Even though a majority of saturnian irregulars lie in the two offset fields (Fig. 1) at any one instant, and the two fields have been observed many times over multiple oppositions, this survey is still somewhat biased towards moons with certain orbits. The lack of coverage north and south of Saturn (see Figure 1) results in moons with $i$ closer to 90 being harder to detect compared to moons with $i$ closer to 0 or 180. This is exemplified by the fact that we never imaged Bestla, the lowest $i$ Norse group member. Another bias, identified in Ashton et al. (2022), is that the glare from Saturn makes it harder to detect moons with smaller $a$ or pericentre. Given these two biases, the Inuit and Gallic subgroups and the low-$i$ Norse moons ($i<$ 151) are likely under-represented in the current sample, and thus have size distributions that are likely more steep in reality than what is seen in Figure 5. The Norse subgroups, on the other hand, are less affected by such observational biases, with the possible exceptions of the low-$a$ part of the Phoebe subgroup and, to a lesser extent, the Kari subgroup. Therefore, we believe that observational biases in our survey should not influence our analysis or conclusions regarding the Mundilfari subgroup.
6 Collisional Families Simulations
Some of the clusters of saturnian irregular moons are tight enough that they could easily be created by a single disruption event (e.g. the Kiviuq subgroup and the Siarnaq subgroup). On the other hand, the Gallic group and the Mundilfari subgroup have comparatively larger orbital spread, and, as such, we want to test the feasibility of a single collision creating either of these two clusters.
We calculate the orbital extent of a debris field by generating $n$ ejecta particles by adding a fixed $\Delta{\bf v}$ in random directions to that of the progenitor velocity at the moment of collision, similar to what was done in Grav et al. (2003). To be judged as to realistically reproducing a clustering, we use the criteria that at least one of the generated ejecta must overlap with the secular ranges of $a$, $e$, and $i$ (found in Jacobson et al. (2022)) of every moon in that cluster. There are aspects which make this current analysis somewhat approximate. Firstly, we did not look at the 3 dimensional aeia-e-i space, just the aea-e and aia-i space. Secondly, when Jacobson et al. (2022) did their analysis, only one of the 64 new moons was announced, so do not have the computed orbital variations. Thus, we cannot know whether the ejecta correctly recreate the full range of either cluster.
6.1 Gallic group
As mentioned in subsection 4.2, the Gallic group could be created by a head-on cratering event near the apex point of Albiorix. Thus, for the orbit of the progenitor moon, we used the averaged $a$, $e$, and $i$ of Albiorix, 0.1092 au, 0.48, and 36.8 respectively. To crudely simulate a head-on cratering collision we restricted the ejecta to having outbound trajectories that are within 90 of the direction of motion of the progenitor. To recreate the comparatively high-$a$ low-$e$ orbit of S/2004 S 24, we set the collision to occur while the progenitor was at apocenter (with true anomaly $f$ = 180). The longitude of ascending node, $\Omega$, argument of periapsis, $\omega$, were set to 0and 90, respectively, to match the small inclination range of the group and the fact that the inclinations of almost all members are greater than that of Albiorix. In order to get ejecta out to the distance of S/2004 S 24 one needs a $\Delta v$ of 330 m/s (see Figure 7). This velocity is significantly higher than what is expected for an irregular moon collisional family (see Turrini et al. (2008) and subsubsection 6.2.1). Additionally, although the ejecta overlap with the $e$ ranges of all Gallic group members (that have ranges), there appears to be a trend (sans S/2004 S 24) that the larger $a$ members also have slightly larger $e$.
We next turn our attention to just the Albiorix subgroup (this is the Gallic group without S/2004 S 24). To get an ejecta cloud that has increasing $e$ with increasing $a$, we changed $\omega$ to 120 and $f$ to 30. To reach the outermost member of the Albiorix subgroup, S/2006 S 12, a $\Delta v$ of 80 m/s is needed (see Figure 7). This much lower $\Delta v$, compared with the Gallic group as a whole, is well within the acceptable range for a collisional family (see Turrini et al. (2008) and subsubsection 6.2.1). Thus, the Albiorix subgroup itself is a good candidates for a collisional family.
As for the origin of S/2004 S 24, perhaps there was a secondary collision where one of the larger $a$ Albiorix subgroup members was impacted while near apocenter and S/2004 S 24 is the largest fragment of that collision. Alternatively, as suggested by Sheppard et al. (2023), S/2004 S 24 might not be related to the rest of the Gallic group and, coincidentally, happen to have similar $i$. Discovering more Gallic group members should lead to a better understanding of the origin of S/2004 S 24.